Solve the equation 3^(5x+2) = 27/ sqrt 3
The real problem I'm having here is understanding how to simplify the fraction with the square root, apparently I can turn 27/ sqrt 3 into 9* sqrt3 somehow, but I don't really understand.

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Solve the equation
3^(5x+2) = 27/ sqrt(3)
-----
3^(5x+2) = 3^3/3^(1/2)
-------
3^(5x+2) = 3^(3-(1/2))
-------
3^(5x+2) = 3^(5/2)
-------
5x+2 = 5/2
5x = 1/2
x = 1/10
---------------
Cheers,
Stan H.
--------------
-----

Answer by solver91311(24713)



The process is called "Rationalizing the Denominator"



You can do this because anything divided by itself is one and anything multiplied by one is unchanged.



John

My calculator said it, I believe it, that settles it